OpenStudy (anonymous):

Between which two integers does /350 lie? 349 and 351 174 and 176 18 and 19 19 and 20

OpenStudy (igreen):

350 - 1 = ? 350 + 1 = ?

OpenStudy (anonymous):



OpenStudy (igreen):

Oh

OpenStudy (igreen):

\(\sf \sqrt{350} \approx 18.71\) So what do you think?

OpenStudy (welshfella):

what are the square numbers either side of 350 ? 18^2 = 19^2 =